OpenStudy (anonymous):

Good quality plywood costs $4.00 per square foot. There is approximately a 15% waste of material due to cutting and fitting. He also figures $200 per month for finishing material, glue, tools, etc. If he charges $10.50 per square foot of finished product, how many square feet of plywood did he buy originally if his profit is $233.40 in one month?

OpenStudy (blacksteel):

Okay, so we know that profit is 233.4, and that profit is 10.5*s - c, where s is square feet done and c is costs. First we need to figure out how much plywood he needs to buy per square foot finished. Since there is a 15% waste, we want to buy enough square feet of plywood x per square foot done to satisfy 0.85x = 1, or x = 1/0.85 = 1.1764 (approximately) Then 233.4 = 10.5s - 4*(1.1764s) - 200. Then we solve for s: s = 74.8 square feet completed. Since the problem is to find how much plywood he bought, we need to then multiply this by our factor 1.1764 to get that 88 square feet of plywood were purchased. (Note that while I used 1.1764 while typing, I used the unrounded result when performing the calculations.)

OpenStudy (anonymous):

So the correct answer is 1.1764??

OpenStudy (anonymous):

whoops.... sorry... so 88 is the correct answer???

OpenStudy (blacksteel):

Yes, 88 is the correct answer

OpenStudy (anonymous):

thanks!!!